The Organizing Committee of the 15-th ILAS Conference thanks

Eunkyung Ahn

Kyungpook National University, Daegu, Korea for delivering the talk

“An extended Lie-Trotter formula and its applications" (with Sejong Kim and Yongdo Lim). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Zeyad Al Zhour

Zarqa Private University, Zarqa, Jordan for delivering the talk

“Matrix Results on Weighted Drazin Inverse and Some

Applications" (with Adem Kilicman). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Milica Andjelic

Center for Research on Optimization and Control, Aveiro, Portugal

for delivering the talk

“One upper bound for the largest eigenvalue of the signless

Laplacian" (with Slobodan Simic). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Marina Arav

Georgia State University, Atlanta, GA, USA for delivering the talk

“Sign Patterns That Require Almost Unique Rank" (with Frank

Hall, Zhongshan Li, Assefa Merid, Yubin Gao). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Antonio Aricò

Dipartimento di Matematica - Universitá di Cagliari, Cagliari, Italy

for delivering the talk

“Signal and Image regularization via antireflective transform". in the Mini-Symposium on Implementation and Application issues in

regularizing least squares and total least squares.

The Organizing Committee of the 15-th ILAS Conference thanks

John Bardsley

University of Montana, Missoula, Montana, USA for delivering the talk

“Truncation Rules for Iterative Deblurring Methods". in the Mini-Symposium on Implementation and Application issues in

regularizing least squares and total least squares.

The Organizing Committee of the 15-th ILAS Conference thanks

Wayne Barrett

Brigham Young University, Provo, Utah for delivering the talk

“Minimum rank of edge subdivisions of complete graphs and wheels" (with Ryan Bowcutt, Mark Cutler, Seth Gibelyou, and

Kayla Owens).

in the Mini-Symposium on Combinatorial Matrix Theory.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

José Barría

Santa Clara University, Santa Clara, U.S.A. for delivering the talk

“The strong closure of the similarity orbit for a class of pairs of

finite rank operators". in the Contributed Talks Sessions.

The Organizing Committee of the 15-th ILAS Conference thanks

Ulrike Baur

Chemnitz University of Technology, Chemnitz, Germany for delivering the talk

“Model Reduction for unstable Systems based on Hierarchical

Matrix Arithmetic".

in the Mini-Symposium on Linear Algebra in Model Reduction.

The Organizing Committee of the 15-th ILAS Conference thanks

Christopher Beattie

Virginia Tech, Blacksburg, VA, USA for delivering the talk

“Interpolatory Projection Methods for Parameterized Model

Reduction".

in the Mini-Symposium on Linear Algebra in Model Reduction.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Tom Bella

University of Connecticut, Storrs, USA for delivering the talk

“Eigenproblems for quasiseparable matrices" (with Yuli

Eidelman, Israel Gohberg and Vadim Olshevsky). in the Mini-Symposium on Eigenproblems: Theory and computation.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Gabriel Bengochea

Universidad Autónoma de la Ciudad de México, México D.F.

for delivering the talk

“Duality in the Hopf Algebra of multivariate polynomials" (with

L. Verde-Star). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Peter Benner

TU Chemnitz, Fakultät für Mathematik, Chemnitz, Germany

for delivering the talk

“Balancing-Related Model Reduction for Large-Scale Unstable

Systems".

in the Mini-Symposium on Linear Algebra in Model Reduction.

The Organizing Committee of the 15-th ILAS Conference thanks

Avi Berman

Technion, Haifa, Israel for delivering the talk

“The Colin de Verdière Parameter- a progress report". in the Mini-Symposium on Combinatorial Matrix Theory.

The Organizing Committee of the 15-th ILAS Conference thanks

Dario Bini

Dipartimento di Matematica, Universitá di Pisa, Pisa, Italy

for delivering the talk

“Fast solution of a certain Riccati Equation through Cauchy-like

matrices" (with Beatrice Meini and Federico Poloni). in the Mini-Symposium on Matrix Functions and Matrix Equations.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Jean-Louis Boimond

LISA - University of Angers, Angers, France for delivering the talk

“On Steady State Controller in Min-Plus Algebra" (with S.

Lahaye). in the Mini-Symposium on Max Algebra.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Gerald Bourgeois

Faculté de Luminy, Marseille, France for delivering the talk

“About the logarithm function over the matrices". in the Contributed Talks Sessions.

The Organizing Committee of the 15-th ILAS Conference thanks

Richard A. Brualdi

University of Wisconsin - Madison, Madison, USA for delivering the talk “A Conjecture in Combinatorial Matrix Theory". in the Mini-Symposium on Combinatorial Matrix Theory.

The Organizing Committee of the 15-th ILAS Conference thanks

Rafael Bru

Univ. Politécnica, Valencia, Spain for delivering the talk

“On some classes of H-matrices" (with Ljiljana Cvetković,

Vladimir Kostić and Francisco Pedroche). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

María Isabel Bueno Cachadina

The University of California at Santa Barbara, Santa Barbara, USA

for delivering the talk

“Algorithms for computing the Geronimus Transformation". in the Contributed Talks Sessions.

The Organizing Committee of the 15-th ILAS Conference thanks

Peter Butkovic

University of Birmingham, Birmingham, United Kingdom

for delivering the talk

“On the permuted max-algebraic eigenvector problem". in the Mini-Symposium on Max Algebra.

The Organizing Committee of the 15-th ILAS Conference thanks

Miguel Carriegos

Universidad de León, León, Spain for delivering the talk “Reachability of regular switched linear systems". in the Contributed Talks Sessions.

The Organizing Committee of the 15-th ILAS Conference thanks

Nieves Castro-González

Universidad Politécnica de Madrid, Madrid, Spain for delivering the talk

“Representations for the generalized Drazin inverse of additive

perturbations" (with M. F. Martínez-Serrano). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Minerva Catral

University of Victoria, Victoria, Canada for delivering the talk

“The Kemeny Constant in Finite Homogeneous Ergodic Markov

Chains".

in the Mini-Symposium on Nonnegative and eventually nonnegative matrices.

The Organizing Committee of the 15-th ILAS Conference thanks

Wei Cheng

National University of Defense Technology, Changsha, P.R. China

for delivering the talk

“One Type of Inverse Eigenvalue Problems in Quaternionic

Quantum Mechanics" (with Liang-gui Feng). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Cristina Corral

Universidad Politécnica de Valencia, Valencia, Spain for delivering the talk

“On Schur complements of H-matrices" (with R. Bru, I.

Giménez, J. Mas). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Vanesa Cortés

Universidad de Zaragoza, Zaragoza, Spain for delivering the talk

“Some properties of the class sign regular matrices and its subclasses" (with J. M. Peña). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Liliana Costa

Centre for Research on Optimization and Control, Aveiro, Portugal

for delivering the talk

“Acyclic Birkhoff Polytope" (with C. M. da Fonseca and Enide

Andrade Martins). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Steve Cox

Rice University, Houston, TX, USA for delivering the talk

“Eigen-reduction of Large Scale Neuronal Networks" (with Tony

Kellems, Derrick Roos and Nan Xiao).

in the Mini-Symposium on Eigenproblems: Theory and computation.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Glória Cravo

University of Madeira and CELC, Funchal, Portugal for delivering the talk

“Controllability of Matrices with Prescribed Blocks". in the Contributed Talks Sessions.

The Organizing Committee of the 15-th ILAS Conference thanks

Henrique F. da Cruz

U.B.I, Covilhã, Portugal for delivering the talk

“On the matrices that preserve the value of the immanant of the

upper triangular matrices" (with Rosário Fernandes). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Tobias Damm

TU Kaiserslautern, Kaiserslautern, Germany for delivering the talk

“Algebraic Gramians and Model Reduction for Different System

Classes".

in the Mini-Symposium on Linear Algebra in Model Reduction.

The Organizing Committee of the 15-th ILAS Conference thanks

Jane Day

San Jose State University, San Jose, CA, USA for delivering the talk “Graph Energy Change Due to Edge Deletion". in the Contributed Talks Sessions.

The Organizing Committee of the 15-th ILAS Conference thanks

Louis Deaett

University of Wisconsin-Madison, Madison, WI, USA for delivering the talk

“The graph and rank of a positive semidefinite matrix". in the Mini-Symposium on Combinatorial Matrix Theory.

The Organizing Committee of the 15-th ILAS Conference thanks

Luz DeAlba

Drake University, Des Moines, USA for delivering the talk

“The Q-matrix completion problem" (with Leslie Hogben and

Bhaba Sharma).

in the Mini-Symposium on Combinatorial Matrix Theory.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Inderjit Dhillon

University of Texas, Austin, USA for delivering the talk

“On some modified root-finding problems" (with Matyas Sustik). in the Mini-Symposium on Eigenproblems: Theory and computation.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Marija Dodig

CELC, Universidade de Lisboa, Lisbon, Portugal for delivering the talk “Singular systems, state feedback problem". in the Contributed Talks Sessions.

The Organizing Committee of the 15-th ILAS Conference thanks

Hamide Dogan-Dunlap

UTEP, El Paso, TX, USA for delivering the talk

“Thinking Modes Revealed on Students’ Responses from an

Assignment on Linear Independence". in the Mini-Symposium on Linear Algebra Education.

The Organizing Committee of the 15-th ILAS Conference thanks

Gregor Dolinar

Faculty of Electrical Engineering, Ljubljana, Slovenia for delivering the talk

“General preservers of quasi-commutativity" (with Bojan

Kuzma). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

María Elena Domínguez

Universidad Politécnica de Madrid, Madrid, Spain for delivering the talk

“General solution of certain matrix equations arising in filter

design applications". in the Contributed Talks Sessions.

The Organizing Committee of the 15-th ILAS Conference thanks

Esther Dopazo

Facultad de Informática. Universidad Politécnica de Madrid, Boadilla del Monte, Madrid, Spain for delivering the talk

“Further results on the representation of the Drazin inverse of a

2 × 2 block matrix" (with M. F. Martínez-Serrano and N.

Castro-González). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Özlem Esen

Anadolu University, Eskisehir, Turkey for delivering the talk “On The Root Clustering of Matrices". in the Contributed Talks Sessions.

The Organizing Committee of the 15-th ILAS Conference thanks

Claudio Estatico

Universitá di Cagliari, Cagliari, Italy for delivering the talk

“Block splitting least square regularization for structured matrices arising in nonlinear microwave imaging"(with J. Nagy,

F. di Benedetto, M. Pastorino, A. Randazzo, and G. Boza). in the Mini-Symposium on Implementation and Application issues in

regularizing least squares and total least squares.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Sherod Eubanks

Washington State University, Pullman, USA for delivering the talk “Generalized Soules Matrices".

in the Mini-Symposium on Nonnegative and eventually nonnegative matrices.

The Organizing Committee of the 15-th ILAS Conference thanks

Heike Fassbender

TU Braunschweig, Braunschweig, Germany for delivering the talk

“On the numerical solution of large-scale sparse discrete-time

Riccati equations" (with Peter Benner).

in the Mini-Symposium on Matrix Functions and Matrix Equations.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Lihong Feng

Faculty of Mathematics, TU Chemnitz, Chemnitz, Germany

for delivering the talk

“Parametric Model Reduction for Systems with Coupled

Parameters" (with Peter Benner).

in the Mini-Symposium on Linear Algebra in Model Reduction.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Rosário Fernandes

Centro de Estruturas Lineares e Combinatórias (CELC), Lisboa, Portugal

for delivering the talk

“Rank partitions and covering numbers under small perturbations of an element". in the Contributed Talks Sessions.

The Organizing Committee of the 15-th ILAS Conference thanks

Josep Ferrer

Universitat Politecnica de Catalunya, Barcelona, Spain for delivering the talk

“Geometric structure of the equivalence classes of a controllable

pair" (with A. Compta and M. Peña). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Carlos Fonseca

Department of Mathematics, University of Coimbra, Coimbra, Portugal

for delivering the talk

“An inequality for the multiplicity of an eigenvalue". in the Contributed Talks Sessions.

The Organizing Committee of the 15-th ILAS Conference thanks

Ajda Fošner

Institute of Mathematics, Physics and Mechanics, Ljubljana, Slovenia

for delivering the talk

“Commutativity preserving maps on real matrices". in the Contributed Talks Sessions.

The Organizing Committee of the 15-th ILAS Conference thanks

Martin Frank

University of Kaiserslautern, Kaiserslautern, Germany for delivering the talk

“An iterative method for transport equations in radiotherapy"

(with Bruno Dubroca). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Roland Freund

University of California, Davis, CA, USA for delivering the talk

“The Effects of Deflation in Projection-Based Order Reduction". in the Mini-Symposium on Linear Algebra in Model Reduction.

The Organizing Committee of the 15-th ILAS Conference thanks

Susana Furtado

Centro de Estruturas Lineares e Combinatórias da U. L, Porto, Portugal

for delivering the talk

“Order Invariant Spectral Properties for Several Matrices" (with

Charles Johnson). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Shigeru Furuichi

Nihon University, Tokyo, Japan for delivering the talk

“On trace inequalities for products of matrices" (with Ken Kuriyama and Kenjiro Yanagi). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Maria T. Gassó

Inst. Mat. Mult. Universidad Politécnica Valencia, Valencia, Spain

for delivering the talk

“The class of Inverse-Positive matrices with checkerboard pattern" (with Manuel F. Abad, and Juan R. Torregrosa).

in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Stephane Gaubert

INRIA and CMAP, Ecole Polytechnique, Palaiseau, France

for delivering the talk

“Using max-plus eigenvalues to bound the roots of a polynomial" (with Marianne Akian (INRIA) and Adrien Brandejsky (ENS

Cachan)). in the Mini-Symposium on Max Algebra.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Martin Gavalec

University of Hradec Králové, Hradec Králové, Czech Republic

for delivering the talk

“Permuted max-min eigenvector problem" (with J. Plavka). in the Mini-Symposium on Max Algebra.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Lev Glebsky

Universidad Autónoma de San Luis Potosi, San Luis Potosi, México

for delivering the talk

“On low rank perturbations of matrices" (with Luis Manuel

Rivera). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Assaf Goldberger

Tel Aviv University, Tel Aviv, Israel for delivering the talk

“An upper bound on the characteristic polynomial of a nonnegative matrix leading to a proof of the Boyle–Handleman

conjecture" (with Michael Neumann). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Jason Grout

Iowa State University, Ames, USA for delivering the talk

“Characterizing graphs with minimum rank at most a given

number over a finite field using polarities of projective

geometries".

in the Mini-Symposium on Combinatorial Matrix Theory.

The Organizing Committee of the 15-th ILAS Conference thanks

Sergey Grudsky

CINVESTAV del I.P.N., México, México for delivering the talk

“Uniform boundedness of Toeplitz Matrices with variable

coefficients". in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Serkan Gugercin

Virginia Tech, Blacksburg, VA, USA for delivering the talk

“A Krylov-Based Descent Approach for the Optimal H_{2} Model

Reduction of Large-Scale Dynamical Systems". in the Mini-Symposium on Linear Algebra in Model Reduction.

The Organizing Committee of the 15-th ILAS Conference thanks

Chun-Hua Guo

University of Regina, Regina, Canada for delivering the talk

“On Newton’s method and Halley’s method for p-th roots of

matrices".

in the Mini-Symposium on Matrix Functions and Matrix Equations.

The Organizing Committee of the 15-th ILAS Conference thanks

José Ramón Guzmán

Instituto de Investigaciones Económicas. UNAM., México, México

for delivering the talk

“Reduction of an Ito’s diffusion input output model for the

determination of square mean stability". in the Contributed Talks Sessions.

The Organizing Committee of the 15-th ILAS Conference thanks

Danny Hershkowitz

Technion, Haifa, Israel for delivering the talk

“On nonnegative sign equivalent and sign similar factorizations of

matrices" (with Allan Pinkus).

in the Mini-Symposium on Combinatorial Matrix Theory.

Luis Verde-Star

Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Iveta Hnětynková

Dep. of Mathematics, Arizona State University, Tempe, Arizona

for delivering the talk

“On solvability of total least squares problem" (with Z. Strakoš

and M. Plešinger).

in the Mini-Symposium on Implementation and Application issues in regularizing least squares and total least squares.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Roger Horn

University of Utah, Salt Lake City, USA for delivering the talk

“A Canonical Form for Quasi-Real Normal Matrices Under Real

Orthogonal Similarity" (with Geoffrey R. Goodson). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Bruno Iannazzo

Dipto. di Fisica e Mat., Università dell’Insubria, Como, Italy

for delivering the talk “Matrix iterations for matrix functions".

in the Mini-Symposium on Matrix Functions and Matrix Equations.

The Organizing Committee of the 15-th ILAS Conference thanks

Bokhee Im

Chonnam National University, Gwangju, Korea(Rep. of) for delivering the talk

“Representations of trilinear products in Comtrans algebras" (with Jonathan D. H. Smith). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Michael Karow

Technische Universitat, Berlin, Germany for delivering the talk

“Pseudospectra and Stability radii for Hamiltonian Matrices". in the Mini-Symposium on Eigenproblems: Theory and computation.

The Organizing Committee of the 15-th ILAS Conference thanks

Steve Kirkland

University of Regina, Regina, Canada for delivering the talk

“Constructing Laplacian Integral Split Graphs" (with N. Abreu,

M. de Freitas and R. Del Vecchio). in the Mini-Symposium on Combinatorial Matrix Theory.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Jacqueline Klasa-Bompoint

Dawson College, Montreal, Canada for delivering the talk

“Few pedagogical scenarios in Linear Algebra with Cabri and

Maple". in the Mini-Symposium on Linear Algebra Education.

The Organizing Committee of the 15-th ILAS Conference thanks

Andre Klein

University of Amsterdam, Amsterdam, The Netherlands for delivering the talk

“Tensor Sylvester matrices and information matrices of multiple

stationary processes".

in the Mini-Symposium on Matrix Functions and Matrix Equations.

The Organizing Committee of the 15-th ILAS Conference thanks

Bhaskara Rao Kopparty

Indiana State University, Terre Haute, IN, USA for delivering the talk “Generalized inverses of infinite matrices". in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Peter Körtesi

University of Miskolc, Miskolc, Hungary for delivering the talk

“Using Linear Algebra in Teaching Hamilton Quaternions and

Graphs". in the Contributed Talks Sessions.

The Organizing Committee of the 15-th ILAS Conference thanks

Thomas Laffey

University College Dublin, Dublin, Ireland for delivering the talk

“Some constructive techniques in the nonnegative inverse

eigenvalue problem" (with Šmigoc, Helena). in the Mini-Symposium on Nonnegative and eventually nonnegative

matrices.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Peter Lancaster

University of Calgary, Calgary, Canada for delivering the talk

“Linearization of Matrix Polynomials" (with A. Amiraslani and

R.W. Corless).

in the Mini-Symposium on Matrix Functions and Matrix Equations.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Gue Myung Lee

Pukyong National University, Busan, Korea for delivering the talk

“Complexity Analysis of the Primal-Dual Interior Point Method for Second-order Cone Optimization Problem" (with Bo Kyung

Choi). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Hosoo Lee

Kyungpook National University, Daegu, Korea for delivering the talk

“Contractions and nonlinear matrix equations on positive definite cones" (with Yongdo Lim). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Steven Leon

University of Massachusetts Dartmouth, Dartmouth, MA

for delivering the talk “When MATLAB Gives "Wrong" Answers". in the Mini-Symposium on Linear Algebra Education.

The Organizing Committee of the 15-th ILAS Conference thanks

Chi-Kwong Li

College of William and Mary, Williamsburg, USA for delivering the talk

“Eigenvalues of the sum of matrices from unitary similarity

orbits" (with Yiu-Tung poon and Nung-Sing Sze). in the Mini-Symposium on Eigenproblems: Theory and computation.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Jean Jacques Loiseau

IRCCyN-CNRS, Nantes, France for delivering the talk

“Robust stability of positive difference equations" (with M. Di

Loreto). in the Mini-Symposium on Max Algebra.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Silvia Machado

Pontificia Universidade Catolica de São Paulo, São Paulo, Brasil

for delivering the talk

“GPEA’s researches about the meta resources in teaching and learning the notion of basis of a vector space" (with Bianchini, B.

L. and Maranhão, M. C. S. A.). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Mirko Maracci

Dept of Mathematics and CSCI, Siena University, Siena, Italy

for delivering the talk

“Basic notions of Vector Space Theory: students’ models and

conceptions". in the Mini-Symposium on Linear Algebra Education.

The Organizing Committee of the 15-th ILAS Conference thanks

Janko Marovt

Institute of Mathematics, Physics and Mechanics, Ljubljana, Slovenia

for delivering the talk

“Homomorphisms of matrix semigroups over division rings from

dimension two to three" (with Gregor Dolinar). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Maria-da-Graça Marques

University of Algarve and CELC, Faro, Portugal for delivering the talk

“Some conditions for the commutativity of matrix patterns"

(with C. R. Johnson). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

José-Javier Martínez

Universidad de Alcalá, Alcalá de Henares, Spain for delivering the talk

“Polynomial regression in the Bernstein basis" (with Ana Marco). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Enide Martins

Centre for Research on Optimization and Control (CEOC), Aveiro,

Portugal for delivering the talk

“On the spectra of some graphs like weighted rooted trees" (with

Rosário Fernandes and Helena Gomes). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

William Martin

North Dakota State University, Fargo, USA for delivering the talk

“Integrating learning theories and application-based modules in teaching linear algebra" (with S. Loch, L. Cooley, M. Meagher,

S. Dexter and D. Vidakovic). in the Mini-Symposium on Linear Algebra Education.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Judith McDonald

Washington State University, Pullman, WA, USA for delivering the talk

“Nonnegative and Eventually Nonnegative Matrices". in the Mini-Symposium on Nonnegative and eventually nonnegative

matrices.

The Organizing Committee of the 15-th ILAS Conference thanks

William McEneaney

University of California, San Diego, La Jolla, United States

for delivering the talk “Max-Plus Bases, Cornices and Pruning". in the Mini-Symposium on Max Algebra.

The Organizing Committee of the 15-th ILAS Conference thanks

Jodi Mead

Boise State University, Boise, USA for delivering the talk

“Calculating Weights in Least Squares Estimation Using the

Chi-squared Method" (with Rosemary Renaut, ASU). in the Mini-Symposium on Implementation and Application issues in

regularizing least squares and total least squares.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Karl Meerbergen

K.U. Leuven, Heverlee, Belgium for delivering the talk

“Recycling Ritz vectors in the parameterized Lanczos method"

(with Zhaojun Bai).

in the Mini-Symposium on Eigenproblems: Theory and computation.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Beatrice Meini

Dipartimento di Matematica, Universitá di Pisa, Pisa, Italy

for delivering the talk

“From algebraic Riccati equations to unilateral quadratic matrix

equations: old and new algorithms" (with Bini, Dario and

Poloni, Federico).

in the Mini-Symposium on Matrix Functions and Matrix Equations.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Hermann Mena

Escuela Politécnica Nacional, Quito, Ecuador for delivering the talk

“Exponential Integrators for Solving Large-Scale Differential

Riccati Equations" (with Benner, Peter).

in the Mini-Symposium on Matrix Functions and Matrix Equations.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Glenn Merlet

CNRS/LIAFA, Paris, France for delivering the talk

“Semi-group of matrices acting on the max-plus projective

space". in the Mini-Symposium on Max Algebra.

The Organizing Committee of the 15-th ILAS Conference thanks

Rana Mikkelson

Iowa State University, Ames, IA, United States of America

for delivering the talk “Minimum Rank of Graphs with Loops". in the Contributed Talks Sessions.

The Organizing Committee of the 15-th ILAS Conference thanks

Thomas Milligan

University of Central Oklahoma, Edmond, OK, USA for delivering the talk

“On Euclidean Squared Distance Matrices" (with Chi-Kwong Li

and Michael Trossett). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Lon Mitchell

Virginia Commonwealth University, Richmond, United States

for delivering the talk

“Orthogonal Removal of Vertices and Minimum Semidefinite Rank" (with Sivaram Narayan). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Julio Moro

Universidad Carlos III de Madrid, Leganés, Spain for delivering the talk

“Structured Holder condition numbers for eigenvalues under fully

nongeneric perturbations" (with María J. Peláez). in the Mini-Symposium on Eigenproblems: Theory and computation.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

DeAnne Morris

Washington State University, Pullman, USA for delivering the talk

“Jordan forms corresponding to nonnegative and eventually

nonnegative matrices" (with McDonald, Judith). in the Mini-Symposium on Nonnegative and eventually nonnegative

matrices.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Ana Moura

Instituto Superior Técnico, UTL, Lisbon, Portugal for delivering the talk

“Skills, Concepts and Models in a Linear Algebra Course" (with

Santos, P. A.). in the Mini-Symposium on Linear Algebra Education.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

James Nagy

Emory University, Atlanta, USA for delivering the talk

“Lanczos Hybrid Regularization for Image Processing Applications" (with Julianne Chung and Dianne O’Leary). in the Mini-Symposium on Implementation and Application issues in

regularizing least squares and total least squares.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Sivaram Narayan

Central Michigan University, Mount Pleasant, Michigan, USA

for delivering the talk

“Linearly Independent Vertices and Minimum Semidefinite

Rank". in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Michael Neumann

Department of Mathematics, University of Connecticut, Storrs, USA

for delivering the talk

“On Optimal Condition Numbers For Markov Chains" (with Sze,

Nung-Sing and Stephen J. Kirkland).

in the Mini-Symposium on Nonnegative and eventually nonnegative matrices.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Dale Olesky

University of Victoria, Victoria, Canada for delivering the talk

“Group Inverses of Matrices with Path Graphs" (with M. Catral

and P. van den Driessche).

in the Mini-Symposium on Combinatorial Matrix Theory.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Vadim Olshevsky

University of Connecticut, Storrs, USA for delivering the talk

“Can One Invert a Matrix via Graph Manipulations?" (with Tom Bella and Pavel Zhlobich). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Vadim Olshevsky

University of Connecticut, Storrs, USA for delivering the talk

“Lipschitz stability of canonical Jordan bases of H-selfadjoint

matrices" (with Tom Bella and Upendra Prasad). in the Mini-Symposium on Eigenproblems: Theory and computation.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Alejandro Palma

Instituto de Física (BUAP), Puebla, México for delivering the talk

“Solution of the linear time-dependent potential by using a solvable Lie algebra^{∗}" (with M. Villa, and L. Sandoval).

in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Marcela Parraguez

Pontificia Universidad Católica de Valparaíso, Valparaíso, Chile

for delivering the talk

“Construction of a vector space schema" (with Oktaç, Asuman). in the Mini-Symposium on Linear Algebra Education.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Pedro Patricio

Departamento de Matemática, Universidade do Minho, Braga, Portugal

for delivering the talk

“Some additive results on Drazin Inverses" (with R. E. Hartwig). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Marta Peña

Universitat Politecnica de Catalunya, Barcelona, Spain for delivering the talk

“Perturbations preserving conditioned invariant subspaces" (with A. Compta and J. Ferrer). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Cecília Perdigão

Faculdade de Ciências e Tecnologia-UNL, Lisboa, Portugal

for delivering the talk

“On the equivalence class graph" (with Rosário Fernandes). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Jan Plavka

Technical University, Kocice, Slovakia for delivering the talk

“On the robustness of matrices in max-min algebra". in the Mini-Symposium on Max Algebra.

The Organizing Committee of the 15-th ILAS Conference thanks

Bor Plestenjak

University of Ljubljana, Ljubljana, Slovenia for delivering the talk

“Numerical methods for two-parameter eigenvalue problems". in the Mini-Symposium on Eigenproblems: Theory and computation.

The Organizing Committee of the 15-th ILAS Conference thanks

Daniela Ponce

University of Hradec Králové, Hradec Králové, Czech Republic

for delivering the talk

“NP-hard problems in extremal algebras tackled by particle

swarm optimization" (with Gavalec, Martin). in the Mini-Symposium on Max Algebra.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

George Poole

East Tennessee State University, Johnson City, USA for delivering the talk “Whatever Happened to Rook’s Pivoting?". in the Contributed Talks Sessions.

The Organizing Committee of the 15-th ILAS Conference thanks

Edgar Possani

ITAM- Instituto Tecnológico Autónomo de México, México, México

for delivering the talk

“Use of models in the teaching of linear algebra" (with Preciado,

G. and Lozano, D.). in the Mini-Symposium on Linear Algebra Education.

Luis Verde-Star Chair of the Organizing Committee

The Organizing Committee of the 15-th ILAS Conference thanks

Volodymyr Prokip

Institute of Appl. Problem for Mech. and Math. NAS of Ukr, L’viv, Ukraine

for delivering the talk

“On the problem of diagonalizability of matrices over a principal ideal domain".

in the Contributed Talks Sessions.

Luis Verde-Star The Organizing Committee of the 15-th ILAS Conference thanks

Vladimir Protasov

Moscow State University, Moscow, Russia for delivering the talk

“p-radii of linear operators and equations of self-similarity". in the Contributed Talks Sessions.

The Organizing Committee of the 15-th ILAS Conference thanks

Rosa E. Pruneda

University of Castilla-La Mancha, Ciudad Real, Spain for delivering the talk

“Complete Orthogonal Decomposition Compared with Direct

Projection Methods" (with Beatriz Lacruz). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Olga Pryporova

Iowa State University, Ames, IA, USA for delivering the talk “Potential Diagonal and D-convergence".

in the Mini-Symposium on Combinatorial Matrix Theory.

The Organizing Committee of the 15-th ILAS Conference thanks

Rosemary A. Renaut

Department of Mathematics and Statistics, Tempe, USA for delivering the talk

“A Newton Iteration for estimating the regularizing parameter for least squares" (with Jodi Mead, Boise State University). in the Mini-Symposium on Implementation and Application issues in

regularizing least squares and total least squares.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Alicia Roca

Dpto. de Matemática Aplicada, Universidad Politécnica, Valencia, Spain

for delivering the talk

“Pencils with Prescribed Constant Subpencils" (with F. C. Silva). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Juan Alberto Rodríguez

Universitat Rovira i Virgili, Tarragona, España for delivering the talk

“The Laplacian Spectrum of Hypergraphs" (with Aida

Kamisalic). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Siegfried M. Rump

Hamburg University of Technology, Hamburg, Germany for delivering the talk

“The ratio between the Toeplitz and the unstructured condition number" (with H. Sekigawa). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Maria Rosaria Russo

Department of Mathematics - University of Padua, Padova, Italy

for delivering the talk

“On some general determinantal identities of Sylvester type" (with Michela Redivo-Zaglia). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Bert W. Rust

National Institute of Standards and Technology, Gaithersburg, MD, USA

for delivering the talk

“A Truncated Singular Component Method for Ill-Posed Problems" (with Dianne P. O’Leary, University of Maryland). in the Mini-Symposium on Implementation and Application issues in

regularizing least squares and total least squares.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Ahmed Salam

Université du Littoral-Côte d’Opale, Calais, France for delivering the talk

“A structure-preserving Arnoldi-like method for a class of

structured matrices". in the Contributed Talks Sessions.

The Organizing Committee of the 15-th ILAS Conference thanks

Salvador Sánchez Perales

Benemérita Universidad Autónoma de Puebla, Puebla, México

for delivering the talk

“Manifold of proper elements" (with S. Djordjevic). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Kenzi Satô

Tamagawa University, Tokyo, Japan for delivering the talk

“The algebraic relations of curvatures of PL manifolds". in the Contributed Talks Sessions.

The Organizing Committee of the 15-th ILAS Conference thanks

Elisa Schaeffer

Universidad Autónoma de Nuevo León, San Nicolás de los Garza, México

for delivering the talk

“Locally computable approximations of absorption times for graph clustering" (with Pekka Orponen and Vanesa Avalos).

in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Burkhard Schaffrin

Ohio State University, Columbus, OH, USA for delivering the talk

“Total least-squares regularization of Tykhonov type and an

ancient racetrack in Corinth" (with Kyle Snow). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Hans Schneider

University of Wisconsin, Madison, Madison, WI, USA for delivering the talk

“Nonnegative linear algebra and max linear algebra: where’s the

difference?". in the Mini-Symposium on Max Algebra.

The Organizing Committee of the 15-th ILAS Conference thanks

Anatoly Sebeldin

University UGANC, Guinea, Conakry, Guinea for delivering the talk

“Algorithm resolving problem of determination of finite cyclicgroup by its automorphism group" (with V. K. Vildanov

and A. L. Sylla). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Morteza Seddighin

Indiana University East, Richmod, Indiana for delivering the talk “Matrix Optimization in Statistics". in the Contributed Talks Sessions.

The Organizing Committee of the 15-th ILAS Conference thanks

Peter Semrl

University of Ljubljana, Ljubljana, Slovenia for delivering the talk “Locally linearly dependent operators". in the Contributed Talks Sessions.

The Organizing Committee of the 15-th ILAS Conference thanks

Hristo Sendov

The University of Western Ontario, London, Canada for delivering the talk

“Spectral Manifolds" (with A. Daniilidis, J. Malick, and A.

Lewis). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Sergey Sergeev

University of Birmingham, Birmingham, United Kingdom

for delivering the talk

“On Kleene stars and intersection of finitely generated

semimodules". in the Mini-Symposium on Max Algebra.

Luis Verde-Star Chair of the Organizing Committee

Bryan Shader

University of Wyoming, Laramie, WY, USA

for delivering the talk

“Average minimum rank of a graph" (with Francesco Barioli, Shaun Fallat, Tracy Hall, Daniel Hershkowitz, Leslie Hogben,

Ryan Martin, and Hein van der Holst).

in the Mini-Symposium on Combinatorial Matrix Theory.

Luis Verde-Star The Organizing Committee of the 15-th ILAS Conference thanks

Mohammad Shahryari

Tabriz University, Tabriz, Iran, Tabriz, Iran for delivering the talk “Z_{2}-graded symmetry classes of tensors". in the Contributed Talks Sessions.

The Organizing Committee of the 15-th ILAS Conference thanks

Ivan Singer

Romanian Academy of Sciences, Bucharest, Romania for delivering the talk “Max-min convexity". in the Mini-Symposium on Max Algebra.

The Organizing Committee of the 15-th ILAS Conference thanks

John Sinkovic

Technische Universiteit Eindhoven, Eindhoven, Netherlands

for delivering the talk

“An upper bound for the maximum nullity of a symmetric matrix whose graph is outerplanar". in the Contributed Talks Sessions.

The Organizing Committee of the 15-th ILAS Conference thanks

Klemen Sivic

Institute of Mathematics, Physics and Mechanics, Ljubljana, Slovenia

for delivering the talk “On varieties of commuting triples". in the Contributed Talks Sessions.

The Organizing Committee of the 15-th ILAS Conference thanks

Helena Šmigoc

University College Dublin, Dublin, Ireland for delivering the talk

“An example of constructing a nonnegative matrix with given

spectrum" (with Laffey, Thomas).

in the Mini-Symposium on Nonnegative and eventually nonnegative matrices.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Graça Soares

University of Tras-os-Montes and Alto Douro, Vila Real, Portugal

for delivering the talk

“Inequalities on an indefinite inner product space" (with N.

Bebiano et al.). in the Contributed Talks Sessions.

The Organizing Committee of the 15-th ILAS Conference thanks

Wolfgang Stefan

Arizona State University, Tempe, Arizona for delivering the talk

“Regularizing Least Squares with the Concentration Method"

(with Rosemary Renaut and Anne Gelb).

in the Mini-Symposium on Implementation and Application issues in regularizing least squares and total least squares.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Marko Stosic

Instituto de Sistemas e Robotica, IST, Lisbon, Portugal for delivering the talk

“On Generalized Procrustes Problem" (with João Xavier). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

David Strong

Pepperdine University, Malibu, California, USA for delivering the talk

“A Java Applet and Introductory Tutorial for the Jacobi,

Gauss-Seidel and SOR Methods". in the Mini-Symposium on Linear Algebra Education.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Jeffrey Stuart

Pacific Lutheran University, Tacoma, USA for delivering the talk “Spectrally Arbitrary Ray Patterns". in the Contributed Talks Sessions.

The Organizing Committee of the 15-th ILAS Conference thanks

Daniel Szyld

Temple University, Philadelphia, USA for delivering the talk

“On General Matrices Having the Perron-Frobenius Property"

(with Elhashash, Abed).

in the Mini-Symposium on Nonnegative and eventually nonnegative matrices.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Daniel Szyld

Temple University, Philadelphia, USA for delivering the talk

“Convergence of Stationary Iterative Methods for Hermitian Semidefinite Linear Systems" (with Frommer, Andreas and

Nabben, Reinhard).

in the Mini-Symposium on Nonnegative and eventually nonnegative matrices.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Bit-Shun Tam

University of Birmingham, Birmingham, UK for delivering the talk

“Maximizing spectral radius of unoriented Laplacian matrix"

(with Ding-Jung Chang and Shui-Hei Wu). in the Mini-Symposium on Combinatorial Matrix Theory.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Denis Tanguay

Université du Québec à Montréal (UQAM), Montréal, Canada

for delivering the talk

“A fundamental paradox in learning algebra" (with Corriveau,

Claudia). in the Mini-Symposium on Linear Algebra Education.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Isabel Teixeira Matos

Centro de Estruturas Lineares e Combinatórias (CELC), Lisboa, Portugal

for delivering the talk

“A Completion Problem over the Field of Real Numbers" (with

Fernando C. Silva). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

María Trigueros

Depto de Matemáticas ITAM, México DF, México for delivering the talk

“Spanning sets and vector spaces they generate: an APOS analysis" (with Ku, Darly and Oktaç, Asuman). in the Mini-Symposium on Linear Algebra Education.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Mitsuru Uchiyama

Shimane University, Matsue, Shimane, Japan for delivering the talk “A New Majorization between functions". in the Contributed Talks Sessions.

The Organizing Committee of the 15-th ILAS Conference thanks

Frank Uhlig

Mathematics, Auburn University, Auburn, AL for delivering the talk

“Convex and Non-convex Optimization Problems for the Field of

Values of a Matrix". in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Ernesto Vallejo

Instituto de Matemáticas, Morelia, México for delivering the talk

“Additivity obstructions for integral matrices and pyramids"

(with Miguel Santoyo). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Pauline van den Driessche

University of Victoria, Victoria, Canada for delivering the talk

“Bounds for the Perron root using max eigenvalues" (with

Elsner, Ludwig). in the Mini-Symposium on Max Algebra.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Hein van der Holst

Eindhoven University of Technology, Eindhoven, The Netherlands

for delivering the talk

“Computing the minimum rank of partial 2-trees".

in the Mini-Symposium on Combinatorial Matrix Theory.

The Organizing Committee of the 15-th ILAS Conference thanks

Kevin Vander Meulen

Redeemer University College, Ancaster, Ontario, Canada for delivering the talk

“Sparse Inertially Arbitrary Sign Patterns" (with L. Vanderspek

and M. Cavers).

in the Mini-Symposium on Combinatorial Matrix Theory.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Paul Van Dooren

Université Catholique de Louvain, Louvain-la-Neuve, Belgium

for delivering the talk

“H2 Approximation and Tangential Rational Interpolation" (with

K. Gallivan and P.A. Absil).

in the Mini-Symposium on Linear Algebra in Model Reduction.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Xaab Nop Vargas

ICYTDF, México, México for delivering the talk

“Students difficulties with the concept of vector space from point of view of APOS Theory". in the Contributed Talks Sessions.

The Organizing Committee of the 15-th ILAS Conference thanks

Luis Verde-Star

UAM Iztapalapa, México DF, México for delivering the talk “Linear algebraic approach to rational functions". in the Contributed Talks Sessions.

The Organizing Committee of the 15-th ILAS Conference thanks

Luis Vieira

Feup, Porto, Portugal for delivering the talk

“Euclidean Jordan algebras and inequalities on the parameters

and on the spectra of a strongly regular graph". in the Contributed Talks Sessions.

The Organizing Committee of the 15-th ILAS Conference thanks

James Weaver

University of West Florida, Pensacola, FL for delivering the talk

“Nonsingularity of Divisor Tournaments" (with Rohan

Hemasinha and Jeffrey L. Stuart). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Piotr Wojciechowski

University of Texas at El Paso, El Paso, TX for delivering the talk

“Orderings of matrix algebras and their applications". in the Contributed Talks Sessions.

The Organizing Committee of the 15-th ILAS Conference thanks

Iwona Wrobel

Warsaw University of Technology and Polish Academy of Sciences,

Warsaw, Poland for delivering the talk

“The Gauss-Lucas theorem and the numerical range". in the Contributed Talks Sessions.

The Organizing Committee of the 15-th ILAS Conference thanks

Pei Yuan Wu

National Chiao Tung University, Hsinchu, Taiwan for delivering the talk

“Numerical ranges of nilpotent operators" (with Hwa-Long Gau). in the Contributed Talks Sessions.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Michelle Zandieh

Arizona State University, Tempe, AZ, USA for delivering the talk

“Design of a unit to teach eigenvectors and eigenvalues based on

the instructional design principles of Realistic Mathematics Education" (with Rasmussen, C. and Larson, C.). in the Mini-Symposium on Linear Algebra Education.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Karel Zimmermann

Faculty of Mathematics and Physics, Charles University, Prague, Czech

Republic for delivering the talk

“Solving two-sided (max,plus)-linear equation systems". in the Mini-Symposium on Max Algebra.

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Juan Carlos Zúñiga

Department of Mathematics, University of Calgary, Calgary, Alberta,

Canada for delivering the talk

“Matrix polynomials, rational matrices and linear systems: A

review" (with Lancaster, Peter).

in the Mini-Symposium on Matrix Functions and Matrix Equations.

Luis Verde-Star Chair of the Organizing Committee